An economy is operating with an output that is $600 billion dollars above its natural rate of $2400 billion dollars and fiscal policymakers want to close the inflationary gap. The central bank agrees to hold the interest rate constant so there is no crowding out. The marginal propensity to consume is 3/4. In which direction and by how much would the government spending need to change to close the gap? Fully explain your answer and provide a graph that shows the initial situation.
Answer #1

Actual GDP = $3,000 billion

Natural rate of GDP = $2,400 billion

Inflationary gap = $600 billion

MPC = (3 / 4) = 0.75

Government spending multiplier = [1 / (1 - MPC)] = [1 / (1 - 0.75)] = 4

To reduce inflationary gap of $600 billion, government spending should reduce by $150 billion which will reduce aggregate demand by $150 billion * Multiplier = $150 billion * 4 = $600 billion.

CRAS Supply Polo ve P Demand (Do? 2900 3000 Puentity